a bomb of 12kg explodes into pieces of masses4kg and 8kg the velocity of 8kg mass is 6m/sthe kinetic energy of other mass is

Grade:12th pass

1 Answers

Arun
6 years ago
By law of conservation of moment conservation:  
initial momentum =final momentum 

12 x 0 = (4xv) + (8 x6) 
So v= 48/4 =12m/s 

Hence the piece having 4kg mass will move with 12m/s in opposite direction of the piece having mass 8kg 

now the KE of 4kg piece= 12x(mv^2) =(1/2)x 4 x 12^2 = 288J